# AnomalyDINO: Boosting Patch-based Few-shot Anomaly Detection with DINOv2

### v1

@@ Abstract
Recent advances in multimodal foundation models have set new standards in few-shot anomaly detection. This paper explores whether high-quality visual features alone are sufficient to rival existing state-of-the-art vision-language models. We affirm this by adapting DINOv2 for one-shot and few-shot anomaly detection, with a focus on industrial applications. We show that this approach does not only rival existing techniques but can even outmatch them in many settings. Our proposed vision-only approach, AnomalyDINO, is based on patch similarities and enables both image-level anomaly prediction and pixel-level anomaly segmentation. The approach is methodologically simple and training-free and, thus, does not require any additional data for fine-tuning or meta-learning. Despite its simplicity, AnomalyDINO achieves state-of-the-art results in one- and few-shot anomaly detection (e.g., pushing the one-shot performance on MVTec-AD from an AUROC of $93.1\%$ to $96.6\%$). The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, for example, in industrial contexts.
@@ 1 Introduction
Anomaly detection (AD) in machine learning attempts to identify instances that deviate substantially from the nominal data distribution $p_{\mathrm{norm}}(x)$. Anomalies, therefore, raise suspicion of being ‘generated by a different mechanism’ (Hawkins, 1980)—often indicating critical, rare, or unforeseen events. The ability to reliably distinguish anomalies from normal samples is highly valuable across various domains, including security (Siddiqui et al., 2019), healthcare (Fernando et al., 2021; Tian et al., 2023), and industrial inspection. In this work, we focus on the latter, where fully automated systems necessitate the ability to detect defective or missing parts to prevent malfunctions in downstream products, raise alerts for potential hazards, or analyze these to optimize production lines. See the right-hand side of Figure 1 for anomalous samples in this context (with predicted anomaly maps from AnomalyDINO).
@@
Figure 1: Anomaly detection with AnomalyDINO based on a single immaculate reference sample (here category ‘Screw’ from MVTec-AD). Whenever suitable we select the relevant patches from the (potentially augmented) reference sample via masking and collect the nominal patch representations in $\mathcal{M}$. For both, masking and feature extraction, we utilize DINOv2. At test time, the distances of patch representations to those in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score after aggregation with $q$. Further examples for other categories are depicted on the right (and in Figures 3 and 4 in Appendix A).
@@
AD for industrial images has gained tremendous interest over the last couple of years. The close-to-optimal results on benchmark data make it seem that the problem of anomaly detection is essentially solved. For instance, Mousakhan et al. (2023) report $99.8\%$ and $98.9\%$ AUROC on the popular benchmarks MVTec-AD (Bergmann et al., 2021) and VisA (Zou et al., 2022), respectively. The most popular AD techniques use the training data to train an anomaly classifier (e.g., Ruff et al., 2020), or a generative model coupled with reconstruction-based (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Mousakhan et al., 2023) or likelihood-based (e.g., Rudolph et al., 2022; Daehyun Kim, 2023) anomaly scoring. However, these approaches operate within the full-shot setting, meaning they rely on access to a sufficiently large amount of training data. Given the challenges associated with dataset acquisition, the attractivity of a fast and easy-to-deploy methodology, and the requirement to rapidly adapt to covariate shifts in the nominal data distribution (Li et al., 2024a), there is an increasing interest in few-shot and zero-shot anomaly detection.
@@
Few-shot techniques, however, heavily rely on meaningful features, or as Reiss et al. (2022) frame it: ‘anomaly detection requires better representations’. Such better representations are now available with the increasing availability and capabilities of foundation models, i.e., large-scale models trained on massive datasets in unsupervised/self-supervised fashion (e.g., Radford et al., 2021; Caron et al., 2021; Oquab et al., 2024). The performance of few-shot anomaly detection techniques has already been boosted by the use of foundation models, mostly by multimodal approaches (e.g., Jeong et al., 2023; Cao et al., 2023; Zhou et al., 2024; Kwak et al., 2024)
@@
Our approach, termed AnomalyDINO, is memory-bank-based and leverages DINOv2 (Oquab et al., 2024) as a backbone (but any architecture with strong patch-level feature extraction is suitable for the proposed approach). We carefully design a suitable preprocessing pipeline for the one-shot scenario, using the segmentation abilities of DINOv2 (which alleviates the additional overhead of another model). At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank.
Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by Chen et al. (2023) or Kwak et al. (2024). Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD (Bergmann et al., 2021) and outperforms all but one competing method on VisA (Zou et al., 2022).
The structure of the paper is as follows: Section 2 reviews relevant prior studies and clarifies the distinctions between the settings addressed by zero- and few-shot, and batched zero-shot techniques. Section 3 introduces our proposed method, AnomalyDINO. An extension of this method to the batched zero-shot scenario is detailed in Appendix D. Section 4 presents the experimental outcomes. Additional results and an ablation study are provided in Appendices A and C, respectively. We address identified failure cases of AnomalyDINO in Appendix B.
@@
  • An extensive analysis demonstrates the efficiency and effectiveness of the proposed approach, outperforming other multimodal few-shot techniques, thereby reducing the gap between few- and full-shot anomaly detection. Specifically, AnomalyDINO achieves state-of-the-art results for few-shot anomaly detection on MVTec-AD, e.g., pushing the one-shot detection from an AUROC of 93.1% to 96.6%. Moreover, our results on VisA are not only competitive with other few-shot methods but also establish new state-of-the-art performance among all training-free few-shot anomaly detectors.
@@ 2 Related Work
Multimodal foundation models have emerged as powerful tools for a wide range of tasks (see e.g., Bommasani et al., 2022; Li et al., 2023; Chen et al., 2020; He et al., 2022; Radford et al., 2021; Liu et al., 2023; OpenAI, 2024). Most relevant to visual AD are multimodal approaches based on CLIP (Radford et al., 2021) or recent LLMs (e.g., OpenAI, 2024), but also vision-only approaches like DINO (Caron et al., 2021; Oquab et al., 2024). CLIP (Radford et al., 2021) learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations. The model uses a contrastive learning objective that aligns the embeddings from image and text encoders, optimizing the similarity between corresponding image-text pairs. This common feature space for vision and language can be utilized for several downstream tasks, such as zero-shot image classification by assessing similarities to a set of class-specific prompts. DINO (Caron et al., 2021; Oquab et al., 2024) leverages a self-supervised student-teacher framework based on vision transformers (Dosovitskiy et al., 2021). It employs a multi-view strategy to train models to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks. DINOv2 (Oquab et al., 2024) combines ideas from DINO with patch-level reconstruction techniques (Zhou et al., 2022) and scales to larger architectures and datasets. The features extracted by DINO are well-suited for anomaly detection as they incorporate both local and global information, are robust to multiple views and crops, and benefit from large-scale pre-training. GroundingDINO (Liu et al., 2023), builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
@@
Given a predefined notion of normality, the anomaly detection task is to detect test samples that deviate from this concept (Yang et al., 2021; Ruff et al., 2021). In this work, we focus on low-level sensory anomalies of industrial image data, i.e., we do not target the detection of semantic anomalies but of low-level features such as scratches of images of industrial products (see e.g., Figure 1). Several works tackled this task by either training an anomaly classifier (e.g., Ruff et al., 2020) or a generative model, which allows for reconstruction-based or likelihood-based AD (e.g., Wolleb et al., 2022; Livernoche et al., 2023; Zhang et al., 2023; Daehyun Kim, 2023; Mousakhan et al., 2023).
@@
However, this work focuses on few-shot AD, which tackles anomaly detection under strong restrictions on the number of available training samples. Various recent methods recognized the effectiveness of pre-trained vision models for few-shot AD. The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets (He et al., 2015), Wide ResNets (Zagoruyko and Komodakis, 2016), and Vision Transformers (ViT, Dosovitskiy et al., 2020). These models can extract representations on a patch and pixel level, which can be compared against a memory bank, which aggregates patch and pixel features provided by the training set. The comparison with the memory bank is typically conducted using a nearest neighbor approach (Bergman et al., 2020; Cohen and Hoshen, 2020; Defard et al., 2020; Roth et al., 2022; Xie et al., 2023; Li et al., 2024b). Loosely speaking, patches and pixels corresponding to anomalies are expected to have a high distance to their nearest neighbor in the memory bank. Another line of work builds upon the success of pre-trained language-vision models in zero-shot classification. The underlying idea consists of two steps. First, these approaches define sets of prompts describing nominal samples and anomalies. Second, the corresponding textual embeddings are compared against the image embeddings (Jeong et al., 2023; Chen et al., 2023; Kwak et al., 2024; Zhou et al., 2024). Images whose visual embedding is close to the textual embedding of a prompt associated with an anomaly are classified as anomalous. However, these methods either require significant prompt engineering (e.g., Chen et al., 2023, use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings). Lastly, another type of few-shot anomaly detection builds upon the success of multimodal chatbots. These methods require more elaborate prompting and techniques for interpreting textual outputs (Xu et al., 2024). Since these methods do not require a memory bank, they are capable of performing zero-shot anomaly detection.
@@
Previous works consider different AD setups, which complicates their evaluation and comparison. To remedy this, we provide a taxonomy of recent few- and zero-shot AD based on the particular ‘shot’-setting, the training requirements, and the modes covered by the underlying models. We categorize three ‘shot’-settings: zero-shot, few-shot, and batched zero-shot. Zero- and few-shot settings are characterized by the number of nominal training samples a method can process, before making predictions on the test samples. In batched zero-shot, inference is not performed sample-wise but based on a whole batch of test samples, usually the full test set. For instance, the method proposed by Li et al. (2024b) benefits from the fact that a significant majority of pixels correspond to normal pixels, which motivates the strategy of matching patches across a batch of images. Another work that considers this setting is that of Li et al. (2024a), who deploy a parameter-free anomaly detector based on the effect of batch normalization. We split the training requirements into the categories ‘Training-Free’, ‘Fine-Tuning’, and ‘Meta-Training’. ‘Training-Free’ approaches do not require any training, while ‘Fine-Tuning’ methods use the few accessible samples to modify the underlying model. In contrast, ‘Meta-Training’ is associated with training the model on a dataset related to the test data. For example, Li et al. (2024a) train their model on MVTec-AD containing all classes except the class they test against. Zhou et al. (2024) and Chen et al. (2023) train their model on VisA when evaluating the test performance on MVTec-AD and vice versa. Finally, we differentiate the leveraged models, which are either vision models (such as pre-trained ViT) or language-vision models (such as CLIP). We provide a detailed summary in Table 1.
Table 1: Taxonomy of recent few- and zero-shot anomaly detection methods. The $\dagger$ indicates approaches that were introduced as full-shot detectors but then considered as few-shot detectors in later works, see e.g., Roth et al. (2022).
Method Setting Training Type Modes
DN2 (Bergman et al., 2020) Few-Shot Training-Free Vision
SPADE (Cohen and Hoshen, 2020) Few-Shot Training-Free Vision
PaDiM (Defard et al., 2020) Few-Shot Fine-Tuning Vision
PatchCore (Roth et al., 2022) Few-Shot Training-Free Vision
PatchCore–opt (Santos et al., 2023) Few-Shot Training-Free Vision
MuSc (Li et al., 2024b) Batched Zero-Shot Training-Free Vision
GraphCore (Xie et al., 2023) Few-Shot Training (GNN) Vision
WinCLIP (Jeong et al., 2023) Zero-/Few-Shot Training-Free Vision + Language
APRIL-GAN (Chen et al., 2023) Zero-/Few-Shot Meta-Training Vision + Language
ADP (Kwak et al., 2024) Few-Shot Fine-Tuning Vision + Language
AnomalyCLIP (Zhou et al., 2024) Zero-Shot Meta-Training Vision + Language
GPT4-V (Xu et al., 2024) Zero-Shot Training-Free Vision + Language
ACR (Li et al., 2024a) Batched Zero-Shot Meta-Training /
AnomalyDINO (Ours) Few-Shot Training-Free Vision
3 Patch-Matching for Visual Anomaly Detection
@@
This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features. The strong feature extraction capabilities of the utilized backbone allow us to fall back on a simple patch-based matching strategy to perform few-shot and training-free AD for image data. We follow a memory-bank approach, i.e., we first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$. Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$. A suitable aggregation of the patch-based distances gives the anomaly score on image-level.
3.1 Anomaly Detection with Nearest Patch Distances

### v2

@@ Abstract
Recent advances in multimodal foundation models have set new standards in few-shot anomaly detection. This paper explores whether high-quality visual features alone are sufficient to rival existing state-of-the-art vision-language models. We affirm this by adapting DINOv2 for one-shot and few-shot anomaly detection, with a focus on industrial applications. We show that this approach does not only rival existing techniques but can even outmatch them in many settings. Our proposed vision-only approach, AnomalyDINO, is based on patch similarities and enables both image-level anomaly prediction and pixel-level anomaly segmentation. The approach is methodologically simple and training-free and, thus, does not require any additional data for fine-tuning or meta-learning. Despite its simplicity, AnomalyDINO achieves state-of-the-art results in one- and few-shot anomaly detection (e.g., pushing the one-shot performance on MVTec-AD from an AUROC of $93.1\%$ to $96.6\%$). The reduced overhead, coupled with its outstanding few-shot performance, makes AnomalyDINO a strong candidate for fast deployment, e.g., in industrial contexts.
@@ 1 Introduction
Anomaly detection (AD) in machine learning attempts to identify instances that deviate substantially from the nominal data distribution $p_{\mathrm{norm}}(x)$. Anomalies, therefore, raise suspicion of being ‘generated by a different mechanism’ [15]—often indicating critical, rare, or unforeseen events. The ability to reliably distinguish anomalies from normal samples is highly valuable across various domains, including security [38], healthcare [13, 39], and industrial inspection. In this work, we focus on the latter, where fully automated systems necessitate the ability to detect defective or missing parts to prevent malfunctions in downstream products, raise alerts for potential hazards, or analyze these to optimize production lines. See the right-hand side of Figure 1 for anomalous samples in this context.
@@
Figure 1: Anomaly detection with AnomalyDINO based on a single immaculate reference sample (here category ‘Screw’ from MVTec-AD). We collect the nominal patch representations from the (potentially augmented) reference sample(s) in the memory bank $\mathcal{M}$. At test time, we select the relevant patch representation via masking (if applicable). The distances of those to the nominal representations in $\mathcal{M}$ give rise to an anomaly map and the corresponding anomaly score $s(\mathbf{x}_{\mathrm{test}})$ using the aggregation statistic $q$. For both, masking and feature extraction, we utilize DINOv2. Further examples for other categories are depicted on the right (and in Figures 4 and 5 in Appendix A).
@@
AD for industrial images has gained tremendous interest over the last couple of years. The close-to-optimal results on benchmark data make it seem that the problem of anomaly detection is essentially solved. For instance, Mousakhan et al. [27] report $99.8\%$ and $98.9\%$ AUROC on the popular benchmarks MVTec-AD [2] and VisA [48], respectively. The most popular AD techniques use the training data to train an anomaly classifier [36], or a generative model coupled with reconstruction-based [40, 25, 27], or likelihood-based [34, 9] anomaly scoring. However, these approaches operate within the full-shot setting, meaning they rely on access to a sufficiently large amount of training data. Given the challenges associated with dataset acquisition, the attractivity of a fast and easy-to-deploy methodology, and the requirement to rapidly adapt to covariate shifts in the nominal data distribution [21], there is an increasing interest in few-shot and zero-shot anomaly detection.
@@
Few-shot techniques, however, heavily rely on meaningful features, or as [32] frame it: ‘anomaly detection requires better representations’. Such better representations are now available with the increasing availability and capabilities of foundation models, i.e., large-scale models trained on massive datasets in unsupervised/self-supervised fashion [30, 5, 29]. The performance of few-shot anomaly detection techniques has already been boosted by the use of foundation models, mostly by multimodal approaches incorporating language and vision[18, 4, 47, 20].
@@
Our approach, termed AnomalyDINO, is memory bank-based and leverages DINOv2 [29] as a backbone (but the approach could be adapted to other architectures with strong patch-level feature extraction). We carefully design a suitable preprocessing pipeline for the one-shot scenario, using the segmentation abilities of DINOv2 (which alleviates the additional overhead of another model). At test time, anomalous samples are detected based on the high distances between their patch representations and the closest counterparts in the nominal memory bank.
Due to its simplicity, AnomalyDINO can be deployed in industrial contexts very easily—in strong contrast to more complex approaches such as those proposed by [7] or [20]. Yet, the proposed method achieves new state-of-the-art performance on anomaly detection in the few-shot regime on MVTec-AD [2] and outperforms all but one competing method on VisA [48].
The structure of the paper is as follows: Section 2 reviews relevant prior studies and clarifies the distinctions between the settings addressed by zero- and few-shot, and batched zero-shot techniques. Section 3 introduces our proposed method, AnomalyDINO. An extension of this method to the batched zero-shot scenario is detailed in Appendix D. Section 4 presents the experimental outcomes. Additional results and an ablation study are provided in Appendices A and C, respectively. We address identified failure cases of AnomalyDINO in Appendix B. The code to reproduce the experiments is available at https://github.com/dammsi/AnomalyDINO.
@@
  • An extensive analysis demonstrates the efficiency and effectiveness of the proposed approach, outperforming other multimodal few-shot techniques in terms of performance and inference speed. Specifically, AnomalyDINO achieves state-of-the-art results for few-shot anomaly detection on MVTec-AD, e.g., pushing the one-shot detection from an AUROC of 93.1% to 96.6% (thereby halving the gap between the few- and full-shot setting). Moreover, our results on VisA are not only competitive with other few-shot methods but also establish a new state-of-the-art for all training-free few-shot anomaly detectors, achieving the best localization performance across all methods.
@@ 2 Related Work
Multimodal foundation models have emerged as powerful tools for a wide range of tasks, see e.g., [3, 22, 6, 16, 30, 24, 28]. Most relevant to visual AD are multimodal approaches based on CLIP [30] or recent LLMs [28], but also vision-only approaches like DINO [5, 29]. CLIP [30] learns visual concepts from natural language descriptions by training on a dataset of images paired with textual annotations. The model uses a contrastive learning objective that aligns the embeddings from image and text encoders, optimizing the similarity between corresponding image-text pairs. This common feature space for vision and language can be utilized for several downstream tasks, such as zero-shot image classification by assessing similarities to a set of class-specific prompts. DINO [5, 29] leverages a self-supervised student-teacher framework based on vision transformers [12]. It employs a multi-view strategy to train models to predict softened teacher outputs, thereby learning robust and high-quality features for downstream tasks. DINOv2 [29] combines ideas from DINO with patch-level reconstruction techniques [46] and scales to larger architectures and datasets. The features extracted by DINO are well-suited for anomaly detection as they incorporate both local and global information, are robust to multiple views and crops, and benefit from large-scale pre-training. GroundingDINO [24], builds upon the DINO framework and focuses on improving the alignment of textual and visual information, enhancing the model’s performance in tasks requiring detailed object localization and multimodal understanding.
@@
Given a predefined notion of normality, the anomaly detection task is to detect test samples that deviate from this concept [43, 35]. In this work, we focus on low-level sensory anomalies of industrial image data, i.e., we do not target the detection of semantic anomalies but of low-level features such as scratches of images of industrial products (see e.g., Figure 1). Several works tackled this task by either training an anomaly classifier [36] or a generative model, which allows for reconstruction-based or likelihood-based AD [40, 25, 45, 9, 27].
@@
However, this work focuses on few-shot AD, which tackles anomaly detection under strong restrictions on the number of available training samples. Various recent methods recognized the effectiveness of pre-trained vision models for few-shot AD. The underlying idea in these approaches is to leverage the feature representations originating from pre-trained ResNets [17], Wide ResNets [44], and Vision Transformers (ViT) [11]. These models can extract representations on a patch and pixel level, which can be compared against a memory bank, which aggregates patch and pixel features provided by the training set. The comparison with the memory bank is typically conducted using a nearest neighbor approach [1, 8, 10, 33, 41, 23]. Loosely speaking, patches and pixels corresponding to anomalies are expected to have a high distance to their nearest neighbor in the memory bank. Another line of work builds upon the success of pre-trained language-vision models in zero-shot classification. The underlying idea consists of two steps. First, these approaches define sets of prompts describing nominal samples and anomalies. Second, the corresponding textual embeddings are compared against the image embeddings [18, 7, 20, 47]. Images whose visual embedding is close to the textual embedding of a prompt associated with an anomaly are classified as anomalous. However, these methods either require significant prompt engineering (e.g., [7] use a total of $35\times 7$ different prompts for describing normal samples) or fine-tuning of the prompt(-embeddings). Lastly, another type of few-shot anomaly detection builds upon the success of multimodal chatbots. These methods require more elaborate prompting and techniques for interpreting textual outputs [42]. Since these methods do not require a memory bank, they are capable of performing zero-shot anomaly detection.
Table 1: Taxonomy of recent few- and zero-shot anomaly detection methods. The $\dagger$ indicates approaches that were introduced as full-shot detectors but then considered as few-shot detectors in later works, see e.g., [33].
Method Setting Training Type Modes
DN2 [1] Few-Shot Training-Free Vision
SPADE [8] Few-Shot Training-Free Vision
PaDiM [10] Few-Shot Fine-Tuning Vision
PatchCore [33] Few-Shot Training-Free Vision
PatchCore–opt [37] Few-Shot Training-Free Vision
MuSc [23] Batched Zero-Shot Training-Free Vision
GraphCore [41] Few-Shot Training (GNN) Vision
WinCLIP [18] Zero-/Few-Shot Training-Free Vision + Language
APRIL-GAN [7] Zero-/Few-Shot Meta-Training Vision + Language
ADP [20] Few-Shot Fine-Tuning Vision + Language
AnomalyCLIP [47] Zero-Shot Meta-Training Vision + Language
GPT4-V [42] Zero-Shot Training-Free Vision + Language
ACR [21] Batched Zero-Shot Meta-Training /
AnomalyDINO (Ours) Few-Shot Training-Free Vision
@@
Previous works consider different AD setups, which complicates their evaluation and comparison. To remedy this, we provide a taxonomy of recent few- and zero-shot AD based on the particular ‘shot’-setting, the training requirements, and the modes covered by the underlying models. We categorize three ‘shot’-settings: zero-shot, few-shot, and batched zero-shot. Zero- and few-shot settings are characterized by the number of nominal training samples a method can process, before making predictions on the test samples. In batched zero-shot, inference is not performed sample-wise but based on a whole batch of test samples, usually the full test set. For instance, the method proposed in [23] benefits from the fact that a significant majority of pixels correspond to normal pixels, which motivates the strategy of matching patches across a batch of images. Another work that considers this setting [21], deploys a parameter-free anomaly detector based on the effect of batch normalization. We split the training requirements into the categories ‘Training-Free’, ‘Fine-Tuning’, and ‘Meta-Training’. ‘Training-Free’ approaches do not require any training, while ‘Fine-Tuning’ methods use the few accessible samples to modify the underlying model. In contrast, ‘Meta-Training’ is associated with training the model on a dataset related to the test data. For example, [21] train their model on MVTec-AD containing all classes except the class they test against. [47] and [7] train their model on VisA when evaluating the test performance on MVTec-AD and vice versa. Finally, we differentiate the leveraged models, which are either vision models (such as pre-trained ViT) or language-vision models (such as CLIP). We provide a detailed summary in Table 1.
3 Matching Patch Representations for Visual Anomaly Detection
@@
This section introduces AnomalyDINO, which leverages DINOv2 to extract meaningful patch-level features. The strong feature extraction capabilities of the utilized backbone allow us to fall back on a simple patch-based matching strategy to perform few-shot and training-free AD for image data. We first gather relevant patch representations of nominal samples in a memory bank $\mathcal{M}$. Then, for each test patch, we compute its distance to the nearest nominal patch in $\mathcal{M}$. A suitable aggregation of the patch-based distances gives the anomaly score on image-level.
Inspired by PatchCore [33], our method simplifies the memory bank framework by selecting DINOv2 as the backbone, thereby reducing the complexity of the feature engineering stage—specifically in deciding which representations or layers to use and how to aggregate them. We also tailor the memory bank concept to the few-shot learning regime by introducing an improved aggregation statistic, and a new preprocessing pipeline that incorporates zero-shot masking and augmentations. The proposed method is described in detail in the following subsections.
3.1 Anomaly Detection via Patch (Dis-) Similarities